THE MICROWAVE SPECTRUM AND STRUCTURE OF HNO_{3}-HCl

\underline{MARK E. OTT}, K.R. LEOPOLD, Department of Chemistry, University of Minnesota, 207 Pleasant St., SE, Minneapolis, MN 55455.

The microwave spectrum of HNO_{3}-HCl has been obtained via pulsed nozzle Fourier transform microwave spectroscopy. The spectrum is consistent with a planer or near-planer structure in which the HNO_{3} hydrogen bonds to the chlorine and the O-H vector is nearly perpendicular to the H-Cl bond. Preliminary analysis indicates a hydrogen bond length of about 2.3 A, which is similiar to that in HCl–HF. The structure is compared with those of HNO_{3}-NH_{3} and HNO_{3}-H_{2}O, both of which have been previously studied in our laboratory.